OpenStudy (anonymous):

Solve:log(x+9)-logx=3 a.0090 b.31303 c.3.2222 d.111

OpenStudy (anonymous):

log(x+9)-log(x)=3 (x+9)/(x)=1000 (x+9)/(1000)=x -999x=-9 x= 1 / 111 x= 0090